# Singular dynamics for discrete weak K.A.M. solutions of exact twist maps

### v1

@@ 4. An explanation of Arnaud’s observation
Similarly, we also have $x_{+}\in\pi_{1}\circ F^{-1}(y,\nabla^{*}u(y))$. By (v) of Proposition 2.2 and item (a) of Remark 2.4, *$\nabla^{*}u(y)\subset\partial\nabla^{+}u(y)$ and then we have that
@@
Since $\nabla^{+}u(y)$ is a closed convex subset of $\mathbb{R}$, its boundary contains at most two different points, that is,
@@
Then noticing $x_{-}<x_{+}$ and using the twist condition of $F$, we claim that
$[x_{-},x_{+}]=\pi_{1}\circ F^{-1}(y,\nabla^{+}u(y)).$
The proof is complete. ∎
Now combining Propositions 3.10 and 4.1, we can manually verify the bijectivity of the commutative diagram below:
{tikzcd}
where $\Gamma^{+}(x,y)=(y,\partial_{2}S(x,y))$, $\Gamma^{-}(x,y)=(x,-\partial_{1}S(x,y))$ and $F$ is the lift of $f$ associated to $S$. We only show that the map $\Gamma^{-}$ in the above diagram is bijective, and the bijectivity of $\Gamma^{+}$ and $F$ could be proved similarly. That is, for any $x\in\mathbb{R}$,
$\Gamma^{-}(x,\Sigma_{+}(x))$ $=(x,-\partial_{1}S(x,\Sigma_{+}(x)))$
$=(x,-\partial_{1}S(x,\pi_{1}\circ F(x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x))))$
$=(x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x)).$
Because the twist condition, which says all $\Gamma^{+},\Gamma^{-}$ and $F$ are $C^{1}$ diffeomorphisms, and the notion of Lipschitz submanifolds is invariant by $C^{1}$ diffeomorphisms, it follows from Proposition 3.9 that all these three graphs $\operatorname{Graph}(\Sigma_{+}),\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u])$ and $\operatorname{Graph}(\nabla^{+}u)$ are Lipschitz submanifolds of $\mathbb{R}^{2}$. This finishes the proof of Theorem 1.3.
5. Singular dynamics on $\operatorname{Graph}(c+\nabla^{+}u_{c})$
In this section, our goal is to analyze the singular dynamics on $\operatorname{Graph}(c+\nabla^{+}u_{c})$ and to provide proofs of Theorems 1.4 and 1.5. Indeed, Theorem 1.4 follows directly from Theorem 1.1. It suffices to show Theorem 1.5.
@@ 5. Singular dynamics on $\operatorname{Graph}(c+\nabla^{+}u_{c})$
Let $u$ be a discrete weak K.A.M. solution. If $u$ is differentiable at $y$, then
$(\Lambda_{0})^{-1}(y,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(y))=F^{-1}(y,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(y)).$
@@
Denote $(x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x)):=F^{-1}(y,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(y))$. Using Lemma 3.2, there exists $x^{\prime}$ such that $(x^{\prime},y)\in\mathcal{NS}(u)$. Then it follows from Remark 2.11 and item (a) of Lemma 2.5 that $\partial_{2}S(x^{\prime},y)=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(y)$. Since $S$ is ferromagnetic, we have $x=x^{\prime}$ and $(x,y)\in\mathcal{NS}(u)$. Hence $\Sigma_{+}(x)=y$. It follows from $y\not\in\operatorname{Sing}(u)$ and Proposition 3.11 that $x=(\Sigma_{+})^{-1}(y)$. Hence $(\Lambda_{0})^{-1}(y,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(y))=(x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x))$. ∎
Notice that once we prove item (ii) of Theorem 1.5, then the item (i) of Theorem 1.5 can be obtained directly from Theorem 2.15.
Proof of (ii) in Theorem 1.5.
Since *$\mathcal{I}^{*}(u)$ is $F$ invariant, for any $(x,p)\in\mathcal{I}^{*}(u)$, we have $(y,p^{\prime}):=F(x,p)\in\mathcal{I}^{*}(u)$. Then $u$ is differentiable at $y$. Using Lemma 5.1, we have
$(x,p)=F^{-1}(y,p^{\prime})=(\Lambda_{0})^{-1}(y,p^{\prime}).$
Hence, $\Lambda_{0}(x,p)=(y,p^{\prime})=F(x,p)$. That is, the restriction of $\Lambda_{0}$ on *$\mathcal{I}^{*}(u)$ equals $F$. ∎
The item (iii) of Theorem 1.5 follows immediately from item (iv) of Theorem 1.4 and the fact $\operatorname{Sing}(u)$ is at most countable (see [CS04, Proposition 4.1.3]).
Proof of (iv) in Theorem 1.5.
Fix $(x,p)\in\operatorname{Graph}(\nabla^{+}u)$. Suppose that for any $n\in\mathbb{Z}_{+}$, we have
$\Lambda_{0}^{n}(x,p)\not\in\mathcal{K}^{*}(u).$
Then by the definition of $\mathcal{K}^{*}(u)$, for any $n\in\mathbb{Z}_{+}$, we have
$\pi_{1}\circ\Lambda_{0}^{n}(x,p)\not\in\operatorname{Sing}(u).$
Using Lemma 5.1, by induction, we have
$\Lambda_{0}^{n}(x,p)=F^{n}(x,p),\quad\forall n\in\mathbb{Z}_{+}.$
Therefore, for any $n\in\mathbb{Z}_{+}$, we have
@@
$F^{n}(x,p)\in\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u).$
That is, $(x,p)\in\mathcal{I}^{*}(u)$. ∎
Appendix
For selfcontainedness, we give proofs for Lemmas 2.5, 2.6 and 2.12.

### v2

@@ 4.2. An explanation of Arnaud’s observation
Similarly, we derive $x_{+}\in\pi_{1}\circ F^{-1}(y,\nabla^{*}u(y))$. Since $u=T^{-}[v]$ is semiconcave, by Proposition A.2 (iv):
$\nabla^{*}u(y)\subset\partial\nabla^{+}u(y).$
Hence:
@@
Since $\nabla^{+}u(y)$ is a closed convex subset of $\mathbb{R}$, its boundary contains at most two different points. Thus:
@@
Observing that $x_{-}<x_{+}$ and using the fact that $p\mapsto\pi_{1}\circ F^{-1}(y,p)$ is a homeomorphism (by Hypothesis 2), we conclude:
$[x_{-},x_{+}]=\pi_{1}\circ F^{-1}(y,\nabla^{+}u(y)).\qed$
Proposition 4.4.
If $u=T^{-}[v]$, $v\in SC(\mathbb{T})$, then the following graphs
$\operatorname{Graph}(\Sigma_{+}[u]),\ \operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u])\ \text{ and }\ \operatorname{Graph}(\nabla^{+}u)$
are topological submanifolds of $\mathbb{R}^{2}$, related by the commutative diagram:
[…]
Here:
  • $\Gamma^{+}(x,y)=(y,\partial_{2}S(x,y))$,
  • $\Gamma^{-}(x,y)=(x,-\partial_{1}S(x,y))$,
  • $F$ is discrete standard map.
Moreover, if $F$ is locally bi-Lipschitz, these three graphs are Lipschitz submanifolds of $\mathbb{R}^{2}$.
Proof.
Step 1: Verifying Bijectivity of the Diagram.
Both $\Gamma^{\pm}$ and $F$ are homeomorphisms on $\mathbb{R}^{2}$, so their restrictions to the relevant graphs are injective. We now prove surjectivity.
(i) Surjectivity of $\Gamma^{-}$: From Proposition 4.1, $T^{+}[u]$ is differentiable everywhere since $u=T^{-}[v]$ with $v\in SC(\mathbb{T})$. By Lemma 2.19 (b), we have:
(6) $-\partial_{1}S(x,\Sigma_{+}[u](x))=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x),\quad\forall x\in\mathbb{R}.$
Thus: $\Gamma^{-}(x,\Sigma_{+}[u](x))=(x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x))$, implying $\Gamma^{-}$ is surjective.
(ii) Surjectivity of $F$: Let $(y,p)\in\operatorname{Graph}(\nabla^{+}u)$. From Lemma 4.3, let
(7) $x:=\pi_{1}\circ F^{-1}(y,p).$
Then $\Sigma_{+}[u](x)=y$. Combining the definition of $F$ and the equation (6), we obtain:
(8) $\pi_{2}\circ F^{-1}(y,p)=-\partial_{1}S(x,y)=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x).$
Hence
$F^{-1}(y,p)=(x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x)),$
proving surjectivity.
Step 2: Lipschitz Regularity of the Graphs.
If $F$ is locally bi-Lipschitz, then $\Gamma^{\pm}$ are locally bi-Lipschitz as well (Remark 1.9). The graph of a Lipschitz continuous function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a Lipschitz submanifold of $\mathbb{R}^{n+1}$. From Proposition 3.5, $\operatorname{Graph}(\Sigma_{+}[u])$ is a $1$-dimensional Lipschitz submanifold of $\mathbb{R}^{2}$. Since the notion of Lipschitz submanifolds is invariant under bi-Lipschitz homeomorphism, we conclude that $\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u])\cap K$ and $\operatorname{Graph}(\nabla^{+}u)\cap K$ are Lipschitz submanifolds for any compact set $K\subset\mathbb{R}^{2}$. By periodicity of $u$ and $T^{+}[u]$, $\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u])$ and $\operatorname{Graph}(\nabla^{+}u)$ are Lipschitz submanifolds globally. ∎
4.3. Singular dynamics on $\operatorname{Graph}(\nabla^{+}u)$
This section analyzes the singular dynamics on $\operatorname{Graph}(\nabla^{+}u)$, culminating in the proofs of Theorems 1.6 and 1.7.
Let $w=T^{-2}[v]$, $u=T^{-}[v]$, and $v\in SC(\mathbb{T})$. Define the forward map:
$\Sigma_{+}[u]\times\Sigma_{+}[w]:\operatorname{Graph}(\Sigma_{+}[u])\rightarrow\operatorname{Graph}(\Sigma_{+}[w])$
given by $(x,\Sigma_{+}[u](x))\mapsto\left(\Sigma_{+}[u](x),\Sigma_{+}[w](\Sigma_{+}[u](x))\right).$ This induces the maps
$\Omega[u]:\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u])\rightarrow\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[w])$
and
$\Lambda[u]:\operatorname{Graph}(\nabla^{+}u)\rightarrow\operatorname{Graph}(\nabla^{+}w)$
illustrated by the commutative diagram:
${{\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u])}}$${{\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[w])}}$${\operatorname{Graph}(\Sigma_{+}[u])}$${\operatorname{Graph}(\Sigma_{+}[w])}$${\operatorname{Graph}(\nabla^{+}u)}$${\operatorname{Graph}(\nabla^{+}w)}$$\scriptstyle{F}$$\scriptstyle{\Omega[u]}$$\scriptstyle{F}$$\scriptstyle{\Gamma^{+}}$$\scriptstyle{\Gamma^{-}}$$\scriptstyle{\Sigma_{+}[u]\times\Sigma_{+}[w]}$$\scriptstyle{\Gamma^{-}}$$\scriptstyle{\Gamma^{+}}$$\scriptstyle{\Lambda[u]}$
@@ 4.3. Singular dynamics on $\operatorname{Graph}(\nabla^{+}u)$
Let $w=T^{-2}[v]$, $u=T^{-}[v]$, and $v\in SC(\mathbb{T})$. Then the map $\Lambda[u]$ satisfies the following:
  1. (i)
    $\Lambda[u](x+1,p)=\Lambda[u](x,p)+(1,0)$.
  2. (ii)
    $\Lambda[u]$ is continuous. If $F$ is locally bi-Lipschitz, then $\Lambda[u]$ is Lipschitz continuous.
  3. (iii)
    $\Lambda[u]$ does not reverse the orientation of the full pseudo-graph.
  4. (iv)
    The image of a vertical segment under $\Lambda[u]$ is a point that belongs to a vertical segment.
@@
For any $(x,p)\in\operatorname{Graph}(\nabla^{+}u)$, we compute as follows:
$\Lambda[u](x,p)$ $=\Gamma^{+}\circ(\Sigma_{+}[u]\times\Sigma_{+}[w])\circ(\Gamma^{+})^{-1}(x,p)$
$=\Gamma^{+}\circ(\Sigma_{+}[u]\times\Sigma_{+}[w])(z,x)\quad\text{ where }(z,x)\text{ satisfies }p=\partial_{2}S(z,x)$
$=\Gamma^{+}(\Sigma_{+}[u](z),\Sigma_{+}[w](x))$
[…]
$=(\Sigma_{+}[w](x),\partial_{2}S(x,\Sigma_{+}[w](x))).$
Since $u=T^{-}[v]$ is semiconcave, the properties of $\Lambda[u]$ follow:
  • Item (i) and the first half of the item (ii) follow directly from Proposition 3.3.
  • Item (iii) follows from Lemma 3.1.
  • The second half of the item (ii), concerning the Lipschitz continuity of $\Lambda[u]$, follows from Proposition 3.5.
  • Item (iv) is a consequence of Proposition 3.7, as it describes how singularities propagate under the map $\Sigma_{+}[u]$.∎
Proposition 4.6.
Let $u$ be a discrete weak K.A.M. solution. Then the map $\Lambda[u]:\operatorname{Graph}(\nabla^{+}u)\righttoleftarrow$ satisfies the following properties:
  1. (i)
    $\Lambda[u](x+1,p)=\Lambda[u](x,p)+(1,0)$.
  2. (ii)
    $\Lambda[u]$ is continuous. If $F$ is locally bi-Lipschitz, then $\Lambda[u]$ is Lipschitz continuous.
  3. (iii)
    $\Lambda[u]$ preserves the orientation of the full pseudo-graph.
  4. (iv)
    The image of a vertical segment under $\Lambda[u]$ is a point that belongs to a vertical segment.
  5. (v)
    $\Lambda[u]$ admits a rotation number for positive iterations:
    $\rho(\Lambda[u]):=\lim_{n\to+\infty}\frac{1}{n}\left(\pi_{1}\circ\Lambda^{n}[u](x,p)-x\right)$
    exists for all $(x,p)\in\operatorname{Graph}(\nabla^{+}u)$ and is independent of $(x,p)$. In fact, $\rho(\Lambda[u])$ is the rotation number of the Aubry set.
Proof.
This proposition follows from Proposition 3.10 and Lemma 4.5. ∎
Let $u$ be a discrete weak K.A.M. solution. Consider the dynamical system $\Lambda[u]:\operatorname{Graph}(\nabla^{+}u)\righttoleftarrow$. For any subset $A\subset\operatorname{Graph}(\nabla^{+}u)$, the $\alpha$-limit set of $A$ is the set of accumulation points of the negative semiorbit of $\Lambda[u]$ started from $A$:
$\alpha(A):=\bigcap_{n\in\mathbb{Z}_{+}}\overline{\{\Lambda^{i}[u](x,p)\mid i\leq-n,(x,p)\in A\}}.$
Proposition 4.7.
@@
Let $u$ be a discrete weak K.A.M. solution. Then the set $\alpha(\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u))$ satisfies the following properties:
  1. (i)
    The set $\alpha(\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u))$ is closed, non-empty and invariant under the dynamics $F$.
  2. (ii)
    the Aubry set $\mathcal{A}^{*}$ can be expressed as
    $\mathcal{A}^{*}=\bigcap_{u}\alpha(\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u)),$
Proof.
Since $u$ is a discrete weak K.A.M. solution, $T^{-}[u]=u+\bar{S}$.
Step 1. Inclusion of $\{(x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x))\mid x\in\mathcal{O}^{n+1}[u]\}\subset\{\Lambda^{i}[u](x,p)\mid i\leq-n,(x,p)\in\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u)\}:$
For any $x\in\mathcal{O}^{n+1}[u]$, by Remark 2.4 (b), $x$ lies in $(\Sigma_{-}[u])^{n}(\mathcal{O}[u])$. Thus, there exists a backward chain $y\in\mathcal{O}[u]$, $y_{-n+1},\dots,y_{-1}$ such that:
$x\in\Sigma_{-}[u](y_{-n+1}),\quad y_{-n+1}\in\Sigma_{-}[u](y_{-n+2}),\quad\cdots\quad y_{-1}\in\Sigma_{-}[u](y).$
Using Lemma 2.5, this chain can be reversed under $\Sigma_{+}[u]$:
$y_{-n+1}\in\Sigma_{+}[u](x),\quad y_{-n+2}\in\Sigma_{+}[u](y_{-n+1}),\quad\cdots\quad y\in\Sigma_{+}[u](y_{-1}),$
leading to:
$(x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x))\in\Lambda^{-n}[u](y,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(y)).$
Step 2. Inclusion of $\{\Lambda^{i}[u](x,p)\mid i\leq-n,(x,p)\in\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u)\}\subset\{(x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x))\mid x\in\mathcal{O}^{n}[u]\}:$
By Proposition 3.7, $u$ remains differentiable along the backward iterates of $\lambda[u]$. If $u$ is differentiable at $y$, $u$ is differentiable on
$\bigcup_{i\leq 0}\Lambda^{i}[u](y,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(y)).$
So if
$(x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x))\in\bigcup_{i\leq-n}\Lambda^{i}[u](y,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(y)),$
the corresponding chain of points $y_{-n+1},\dots,y_{-1}$ satisfy
$y_{-n+1}\in\Sigma_{+}[u](x),\quad y_{-n+2}\in\Sigma_{+}[u](y_{-n+1}),\quad\cdots\quad y\in\Sigma_{+}[u](y_{-1}),$
and $u$ is differentiable at $y_{-n+1},\dots,y_{-1}$. By Lemma 3.4 and Proposition 3.6, we have
$x\in\Sigma_{-}[u](y_{-n+1}),\quad y_{-n+1}\in\Sigma_{-}[u](y_{-n+2}),\quad\cdots\quad y_{-1}\in\Sigma_{-}[u](y).$
This implies $x\in\mathcal{O}^{n}[u]$.
Step 3. Equivalence of $\alpha(\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u))$ and $\{(x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x))\mid x\in\mathcal{O}^{\infty}[u]\}$:
Using the inclusions from Steps 1 and 2, we deduce:
$\bigcap_{n\in\mathbb{Z}_{+}}\{(x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x))\mid x\in\mathcal{O}^{n}[u]\}=\bigcap_{n\in\mathbb{Z}_{+}}\overline{\{\Lambda^{i}[u](x,p)\mid i\leq-n,(x,p)\in\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u)\}}.$
Thus,
$\alpha(\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u))=\{(x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x))\mid x\in\mathcal{O}^{\infty}[u]\},$
showing by Proposition 2.24 it is closed, non-empty and invariant. By Proposition 2.17,
$\bigcap_{u\text{ weak K.A.M.}}\alpha(\operatorname{Graph}(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u))=\bigcap_{u\text{ weak K.A.M.}}\left\{(x,p)\mid x\in\mathcal{O}^{\infty}[u],p=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x)\right\}=\mathcal{A}^{*}.$
This completes the proof of both parts of the proposition. ∎
Appendix A Semiconcave functions
Let us recall some fundamental definitions and properties on semiconcave functions (see [CS04, Ber07] for instance).
Definition A.1.
Let $u:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a continuous function.
  • For any $x\in\mathbb{R}^{d}$, the sets
  • Let $\Omega\subset\mathbb{R}^{d}$ be a convex open set. We say that $u$ is semiconcave (with a linear modulus) on $\Omega$ if there exists a constant $C\geq 0$ such that
[…]
  • For any $x\in\mathbb{R}^{d}$, a covector $p\in(\mathbb{R}^{d})^{*}$ is called a reachable differential of $u$ at $x$ if there exists a sequence $\{x_{n}\}\subset\mathbb{R}^{d}\backslash\{x\}$ such that $u$ is differentiable at $x_{k}$ for each $k\in\mathbb{Z}_{+}$, and
    $\lim_{k\rightarrow+\infty}x_{k}=x,\quad\lim_{k\rightarrow+\infty}\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u\left(x_{k}\right)=p.$
    The set of all reachable gradients of $u$ at $x$ is denoted by $\nabla^{*}u(x)$.
One can then have the following well-known facts.
Proposition A.2.
Let $u:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a continuous function, $\Omega\subset\mathbb{R}^{d}$ be a convex open set and $x,x_{0}\in\Omega$.
  1. (i)
    Let $u:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a semiconcave function on $\Omega$ with a linear modulus and some semiconcave constant $C$. Then
    $u(x)-u(x_{0})\leq p_{0}(x-x_{0})+\frac{C}{2}\|x-x_{0}\|^{2}$
    for any $p_{0}\in\nabla^{+}u(x_{0})$. Conversely, if there exists $C\geq 0$ and $p_{0}\in(\mathbb{R}^{d})^{*}$ such that
    $u(x)-u(x_{0})\leq p_{0}(x-x_{0})+\frac{C}{2}\|x-x_{0}\|^{2},\ \quad\forall x\in\Omega$
    then $u$ is semiconcave with semiconcavity constant $C$ and $p_{0}\in\nabla^{+}u(x_{0})$.
  2. (ii)
    Let $v:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be another continuous function. If $u\leq v$ on $\Omega$ and $u(x)=v(x)$, then $\nabla^{+}u(x)\supset\nabla^{+}v(x)$ and $\nabla^{-}u(x)\subset\nabla^{-}v(x)$.
  3. (iii)
    $u$ is differentiable at $x$ if and only if $\nabla^{+}u(x)$ and $\nabla^{-}u(x)$ are both nonempty; in this case, we have that
  4. (iv)
    Let $u:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a semiconcave function on $\Omega$. Then $\nabla^{*}u(x)\subset\partial\nabla^{+}u(x)$ for all $x\in\Omega$, where $\partial$ is the boundary taken with respect to the standard topology of $\mathbb{R}^{d}$.
  5. (v)
    Let $u:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a semiconcave function on a compact set $\Omega$. Then the gradient of $u$, defined almost everywhere in $\Omega$, has bounded variation. Hence the gradient of $u$ is bounded in $\Omega$.
Appendix B Twist maps of the $2$-dimensional annulus
We begin by introducing the notion of twist maps.
Definition B.1.
  1. (1)
    $f$ is isotopic to identity ;
  2. (2)
  3. (3)
    $f$ twists verticals to the right: if $F:\mathbb{R}\times\mathbb{R}^{*}\righttoleftarrow$, denoted by $F(x,p)=(y,p^{\prime})$, is a lift of $f$, then there exists $\varepsilon>0$ such that
Definition B.2.
  1. (1)
    $S$ is diagonal-periodic:
    $\forall(x,y)\in\mathbb{R}^{2},\quad S(x+1,y+1)=S(x,y);$
  2. (2)
    there exists $\varepsilon>0$ such that
    $\frac{\partial^{2}S}{\partial x\partial y}(x,y)<-\varepsilon,\quad\forall(x,y)\in\mathbb{R}\times\mathbb{R}.$
  3. (3)
Remark B.3.
For any lift $F$ of a twist map $f$, there exists a unique generating function $S$, up to addition by a constant, that satisfies Definition B.2 (see [MMS89, Appendix 1]).
Proposition B.4.
Let $f$ be a twist map and $S$ be a generating function. For any $c\in\mathbb{R}$, the function
$S_{c}(x,y):=S(x,y)+c(x-y)$
satisfies Hypothesis 1, Hypothesis 2 and Hypothesis 3.
Proof.
The diagonal-periodicity and ferromagnetic property of $S_{c}$ follows directly from items (1) and (3) of Definition B.2, respectively. Since $S$ is of class $C^{2}$, it follows that $S_{c}$ is of class $C^{1}$ and locally semiconcave (see [CS04, Proposition 2.1.2]). The remaining task is to demonstrate that $S_{c}$ is coercive. This property follows from the superlinearity of the generating function $S$, expressed as
$\lim_{|x-y|\rightarrow+\infty}\frac{S(x,y)}{|x-y|}=+\infty.$
This conclusion is an immediate consequence of a lemma from the work of MacKay, Meiss and Stark [MMS89, Lemma A1.6]. Furthermore, the Non-crossing Lemma [ALD83] can be obtained by integrating $\frac{\partial^{2}S}{\partial x\partial y}$ over a quadrangle; we omit its detailed proof here. ∎
Remark B.5.
  1. (a)
    The function $S_{c}$ is referred to as the generating function at the cohomology class $c$ of $\mathbb{T}$. As proved in Proposition B.4, all theories in this paper are applicable to $S_{c}$. For instance, consider any $c$ in the first cohomology group $H^{1}(\mathbb{T},\mathbb{R})$, which is isomorphic to $\mathbb{R}$, the Discrete Weak K.A.M. Theorem 2.8 defines the effective interaction $\overline{S_{c}}$. The function $\alpha(c)=-\overline{S_{c}}$ is called Mather’s $\alpha$-function, which is $C^{1}$. As a second example, the Aubry set $\mathcal{A}_{c}^{*}$ at cohomology class $c$ is $F_{c}$-invariant, where
    $F_{c}(x,p)=F(x,p+c)-(0,c).$
  2. (b)
[…]
Appendix C Proofs of Lemmas 2.19 and 2.22
For selfcontainedness, we give proofs for Lemmas 2.19 and 2.22.